OpenStudy (tamara7301):

equation is y=-2x+2 what is the slope plz show how u found it m=y2-y1 x2-x1

OpenStudy (jango_in_dtown):

y=mx+c has the slope m here we have y=-2x+2 so comparing with the slope intercept form we have m=-2 hence the slop is -2

OpenStudy (texaschic101):

you would use the slope formula : (y2 - y1) / (x2 - x1) if you had 2 sets of points. But when you just have an equation, all you have to do is put the equation in y = mx + b form, where m is your slope. y = -2x + 2 -- it is already in y = mx + b form....the number in the m position, the slope, is -2. questions ?

OpenStudy (tamara7301):

oh thank u so much that helps alot

OpenStudy (texaschic101):

sure thing :)